A particular type of bacteria grows at a rate of 49 % per day. If my dog got infected with 4000 bacterium when her leg was cut by barbed wire, how many bacterium will be present in 24 days when I finally decide I need to take the dog to the veterinarian.
___________bacterium
Help

Answer by jim_thompson5910(35256)

A = P(1+r)^t

A = 4000(1+0.49)^t

A = 4000(1.49)^t

A = 4000(1.49)^24

A = 4000(14,337.4013230579)

A = 57,349,605.2922316

So that's the rough number of bacterium that will be present in 24 days